OpenStudy (anonymous):

Sec(2θ+10°)=csc(4θ+20°) do i set this equal to 180 or 90? the question says to assume that there all acute angles.

OpenStudy (anonymous):

For everybody whos help me at this moment, i love you.

OpenStudy (anonymous):

\[1\div(\cos (2*\theta+10))=1\div(\sin(4*\theta+20))\] \[\cos(2*\theta+10)=\sin(4*\theta+20)\] If we let quantity (2*theta+10) equal to a variable u, then we can see \[\cos (u)=\sin(2*u)\] What angle is so that the sin of its double is it's cosine? I'll give you a hint, it rhymes with "birdy."

OpenStudy (anonymous):

You sir are amazing, i wish you the best life.

OpenStudy (anonymous):

Thank you kind fellow, just came here asking for some help myself, thought I'd pay it forward a little :)